how many ounces of pure water must be added to 60oz of a solution that is 20% water to make a solution that is 30% water?
Answer by jorel1380(3719)

.20(60)+x=.30(60+x)
12+x=18+.30x
.70x=6
x=6/.70=60/7=8 4/7 oz..